OpenStudy (anonymous):

[Part 1 (4 points)] Show all work in dividing 4 t squared minus 1 all over 15 t plus 10 divided by 4 t squared minus 4 t plus 1 all over 3 t squared plus 5 t plus 2. [Part 2 (2 points)] Use complete sentences to explain your process.

OpenStudy (campbell_st):

try factoring the quadratics \[\frac{4t^2 -1}{15t + 10} \div \frac{4t^2 + 4t + 1}{3t^2+5x + 2}\] startinf witht he 1st fraction you get \[\frac{(2t -1)(2t + 2)}{5(3t + 2)} \] factor the 2nd fraction and remember, when dividing by a fraction flip and multiply the 2nd fraction.

OpenStudy (anonymous):

haha see i don't know exactly how to do this could you help?

OpenStudy (campbell_st):

well the 2nd fraction is \[\frac{(2t +1)^2}{(3t + 2)(t + 1)}\] so you problem is now \[\frac{(2t -1)(2t + 1)}{5(3t + 2)} \times \frac{(3t +2)(t + 1)}{(2t + 1)^2}\] remove the common factors then complete the multiplication for the answer.

OpenStudy (anonymous):

Yea i'm trying to do this i just don't know how to start it can you like help me with the answer and ill get another problem and see if i can solve it the way you taught me after this one?